A step down transformer has 3200 windings in the primary coil and a primary voltage of 100 V. determine the number of windings in the secondary coil if the secondary voltage is 45 V.


1
Expert's answer
2021-01-26T17:48:53-0500

By the ideal transformer equation, we have:


"\dfrac{V_S}{V_P}=\dfrac{N_S}{N_P},""N_S=N_P\dfrac{V_S}{V_P},""N_S=3200\ windings\cdot\dfrac{45\ V}{100\ V}=1440\ windings."